OpenStudy (anonymous):

medal and fan! How many solutions can be found for the equation 4x = 4x?

Mehek (mehek14):

infinite any number you put for x will always be true on both sides

OpenStudy (anonymous):

ohh i understand if i have another questin can i ask you its on the same kind of subject @Mehek14

Mehek (mehek14):

go ahead

OpenStudy (anonymous):

How many solutions can be found for the equation 3y + 5 − 2y = 11? @Mehek14

OpenStudy (anonymous):

@Vocaloid @Peaches15

Vocaloid (vocaloid):

well, try solving it and you'll see

Vocaloid (vocaloid):

3y - 2y = __y fill in the blank

OpenStudy (anonymous):

11?

Vocaloid (vocaloid):

3-2 = ?

OpenStudy (anonymous):

How many solutions can be found for the equation 3y + 5 − 2y = 11? Zero One Two Infinitely many

Vocaloid (vocaloid):

3-2 = ?

OpenStudy (anonymous):

one?

Vocaloid (vocaloid):

right, so 3y-2y = ?

OpenStudy (anonymous):

only one solution>11?

OpenStudy (anonymous):

or one

Vocaloid (vocaloid):

3-2 = 1 3y-2y = __y fill in the blank, it follows the same pattern...

OpenStudy (anonymous):

1y?

Vocaloid (vocaloid):

right, 1y, or just y so we have y + 5 = 11 so what is y?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

infinity mano solutions?

Vocaloid (vocaloid):

y = 6 so there's only ONE solution

Vocaloid (vocaloid):

there is only one value (6) that will make the equation true...

OpenStudy (anonymous):

There is only one solution since... 3y+5−2y=11 SSimplify both sides of the equation. 3y+5−2y=11 3y+5+−2y=11 (3y+−2y)+(5)=11(Combine Like Terms) y+5=11 y+5=11 Subtract 5 from both sides. y+5−5=11−5 y=6

OpenStudy (anonymous):

ohh thats a great explanation thanks so much @ex

imqwerty (imqwerty):

Sry i though u guys were talkin abt the original question

OpenStudy (anonymous):

oh its fine thanks anyways

OpenStudy (anonymous):

Welcome.